OpenStudy (anonymous):

Solve this system. Help please 4x-2y=12 3y-4x=1 Answer ( , )

jimthompson5910 (jim_thompson5910):

3y-4x=1 is the same as -4x+3y=1

jimthompson5910 (jim_thompson5910):

add the equations to get 4x-2y=12 -4x+3y=1 ---------- 0x+1y = 13 so 1y = 13 or y = 13

jimthompson5910 (jim_thompson5910):

Use this value of y to get the value of x. I'll let you do this.

OpenStudy (ybarrap):

Or just graph y = 2x/3 + 6 and y=4x/3 + 1/3 and see where they intersect. These are the same equations you have above but I've isolated the y for you. Then set each equal to each other and solve for x like this: 2x/3 + 6 = 4x/3 + 1/3. Now all you need to do is solve for x and then plug into any equation to get y.